Linda R. Coney – 24th April 2009 MOM Update Linda R. Coney 21 September, 2009.
Published byModified over 5 years ago
Presentation on theme: "Linda R. Coney – 24th April 2009 MOM Update Linda R. Coney 21 September, 2009."— Presentation transcript:
Linda R. Coney – 24th April 2009 MOM Update Linda R. Coney 21 September, 2009
Linda R. Coney – 21 Sept 2009 Outline Target Taking Data/Operations Documentation Electrical Continued Run Plan …..
Linda R. Coney – 21 Sept 2009 Target Operations & Data Taking Ran 8 hours Wednesday Used 3,200 target pulses Target operation u First operation without Sheffield experts in person (although one on phone for Target DAQ nuances) u Worked well – no problems u ISIS allowing 1V losses u Dipped target into beam until ~700mV losses s Losses drifted down to ~600mV over the evening Working on DAQ issues u Running at higher particle rate – DS operational, high loss levels in ISIS u Necessitates changes in DAQ to accommodate high particle rate Pion beam at 330 MeV/c – Decay Solenoid beam dynamics study u Nominal setting at 550 A (3.1 T) u 9 settings higher and lower in steps of 0.15 T u Only required ~60 target pulses/point to get sufficient counts in TOF0
Linda R. Coney – 21 Sept 2009 Target Operations Target I stability from 5 Sept 2009 All plots to follow show characteristic two peaks (exact capture position). The two peaks are always the same distance apart - 0.13 mm - corresponding to the resolution on the optical readout of 0.15 mm. The widths of the peaks has not changed since the start of running in ISIS on 5/9/09
Linda R. Coney – 21 Sept 2009 Target Operations II Target I stability from 12 Sept 2009
Linda R. Coney – 21 Sept 2009 Target Operations III Target I stability from 16 Sept 2009 deltaD for 5 th =.13 and for 16 th =.12 Running at same depth – consistent behavior
Linda R. Coney – 21 Sept 2009 Target Operations IV Beam Center Distance over Evening shift Wednesday Shows the time series for the last run, patchy because we were stopping and starting the target a lot. The important thing is the BCD consistency.
Linda R. Coney – 21 Sept 2009 Recent Data See Marco’s slides
Linda R. Coney – 21 Sept 2009 Operations I Updated documentation - DONE Some trips of beamline magnets u usually trip on power supply water flow DAQ u having some trouble with new fADCs u Possible timing issue with TDC and fACD signals u Possible fix this week “invalid” alarms constantly going off on alarm handler extremely taxing/annoying u Problem solved!! Require alarm to be valid for longer than 1 second before audible alert. u No invalid alarms, and significantly reduced number of alarms on water flow. u Underlying communication problem causing invalid alarms still under investigation.
Linda R. Coney – 21 Sept 2009 Operations II Electrical u Sent list of typical magnet current combinations to Daresbury u Calculations being done to determine if running near limits of distribution box on wall Shifters u New list of instructions for novice shifters – Paul Harrison will upload to MICE shifter schedule website u New shifter folder with operating instructions and safety class requirements u Shifters need to be prepared before arriving in MICE LCR for shift! s Big responsibility, full time attention required, mistakes could be critical eLOG u Still hanging and making operations difficult u Need it to work smoothly u Moving to heplnw17 this week – mirror on MICE website from IIT Slow ramp on beamline magnets u In progress
Linda R. Coney – 21 Sept 2009 Next Run Plan for this week: u Commission Decay Solenoid – DONE s Pion optics - Scan through decay solenoid current settings s Compare beam profile in detectors to simulation u Commission FNAL BeamProfileDetectors - DONE u Commission DAQ s Adjust to higher particle rate s Change TOF0 trigger to AND of two planes instead of OR s Set up TOF1 trigger – necessary due to increased particle rate s Figure out why TDC and fADC information for particles not saved appropriately s Commission TOF1 and KL DAQ s Check with Analysis code u Calibrate TOF system s 330 MeV/c pion beam, 150 MeV/c e+ beam, 100 MeV/c e+ beam u Optimize Upstream Beamline s Marco and Chris Rogers s Optimize Q1-Q3 s Move quads to maximize rates in detectors
Linda R. Coney – 21 Sept 2009 Target Operations The plot shows the Beam Centre Distance of the target over time for the entire Saturday shift. The double peak effect can be clearly seen but the important point is that the target was absolutely stable over this prolonged run period. Also the higher points at the beginning and end of the run are when the target was lowered and raised respectively.